OpenStudy (anonymous):

A factory kept on increasing its profit, the same percent every year. Find the percent if it is known that the output doubled in last 2 years.

OpenStudy (dumbcow):

Let P be original profit, r be rate increase each year P(2) = P(1+r)^2 = 2P --> (1+r)^2 = 2 1+r = sqrt2 r = sqrt2 - 1 r = .414 or 41.4%